Diane dives in a pool moving at 7.6m/s until reaching 8.42m from the diving board. Solve the following

A. The average acceleration of the diver as she travels through the water
B. Time the divers come to a stop
C. The speed of the diver after diving 5.87m
1
Expert's answer
2021-10-11T08:55:53-0400

A The average acceleration is


"a=\frac{v^ 2}{2d}=3.44\text{ m/s}^2."

B. The time is


"T=\frac{v}{a}=2.21\text{ s}."

C. The speed after 5.87 m is


"v=\sqrt{2ad}=6.35\text{ m/s}."